OpenStudy (anonymous):

square root of a-6=3

OpenStudy (anonymous):

\[\sqrt {a-6} = 3\] it looks like this, isn't it?

OpenStudy (anonymous):

\[\sqrt{a-6} = 3\] \[a-6 =3^2\] \[a-6 = 9\] \[ a = 9+ 6\] \[a = 15\]